# TL;DR: Too Long, Do Re-weighting for Effcient LLM Reasoning Compression

### v1

@@ 1 Introduction
Direct approaches aimed at addressing the issues of cognitive redundancy and excessive deliberation within reasoning LLMs. Training-free methods CoD ; yao2025activation ; TALE-EP include some that control the internal states of the model during reasoning through prompts or confidence-based techniques to compress the model. Alternatively, the pathway, exemplified by model merging, involves intervening in the parameters of the reasoning LLM to produce relatively concise solutions. Training-based methods primarily focus on sampling and synthesizing relatively concise reasoning paths on specified problem sets through various strategies TokenSkip ; TOPS ; CoT-Valve . These methods involve performing reinforcement learning SimPO ; thinkpruner ; O1-Pruner ; L1 or supervised fine-tuning (SFT) Overthink on reasoning LLMs, enabling the model to learn to generate more concise yet still correct reasoning paths.
@@
We base our approach on an intuitive motivation: when a model is thinking too long, it should reweight more intuitive reasoning paths to simplify the thinking process. Conversely, when the thinking is too direct, it should incorporate more slow-thinking reasoning chains to encourage deeper contemplation. We propose a dynamic Thinking Length Data Re-Weighting method (TlDr), which dynamically balances the model’s complex reasoning using long CoT and efficient reasoning using short CoT data, enabling the model to eliminate redundant cognitive processes. First, we construct short CoT data for simple problems and long CoT data for complex problems. The model begins with an initial ratio and performs reasoning compression using mixed data. After completing a compression cycle, the model re-evaluates the expected benefits of System-1 CoT data and System-2 CoT data to achieve improved performance. Specifically, and in line with intuition, System-1 CoT data can enhance efficiency, so we use an efficiency metric to measure the expected benefit of System-1 data. System-2 CoT, on the other hand, improves reasoning accuracy, and we use an accuracy metric to measure the benefit of System-2 data in terms of reasoning capability.
Compared to various methods requiring fine-tuning data with different reasoning lengths, our approach enables dynamic ratio learning by utilizing the self-sampled LongCoT model and the short CoT data constructed by the original instruct/base model. Through experiments on DeepSeek-Distill-7B/14B, our model achieves excellent compression results on the 7B/14B models, with only a slight decrease in reasoning capability.
@@ 5 Related Work
Despite the strong generalization and reasoning abilities demonstrated by the system 2 reasoning paradigm, the auto-regressive nature of LLMs imposes a significant reasoning burden, limiting their application in agent-based or edge scenarios OS-ATLAS . To address this, various approaches have emerged to improve the reasoning efficiency. These methods can be broadly categorized into two types. One category focuses on building adaptive reasoning-budget. Within this, some training-free methods like CoD CoD and TALE-EP TALE-EP impose budget constraints to control overall reasoning cost. Budget-sensitive models such as L1 L1 , TOPS TOPS , o1 Pruner o1_pruner , and K1.5 team2025kimi add length penalties during the post-training. Some work CoT-Valve ; jiang2025drp ; chain-of-reasoning synthesizes diverse-length CoT data, while TOPS TOPS samples budget-sensitive versions using a data model, and C3oT C3oT compresses original LLM output to train jointly on shorter CoTs. Other approaches involve external models or switching mechanisms Reasoning-in-Flux to allocate budgets. Routellm RouteLLM uses multiple routers to find the most suitable reasoning model, while Self-REF Self-REF employs confidence scores to route based on reasoning difficulty. Another category concentrates on building efficient representations. TokenSkip TokenSkip selects data based on token importance for compressed reasoning and more concise thought chains. COCONUT COCONUT explores more efficient reasoning in the latent space. ICoT-KD CCoT and CCoT CCoT attempt to build more efficient reasoning strategies in the hidden space, while Token Assorted combines hidden space and text-based reasoning to balance interpretability and efficiency. Heima Heima extends hidden-space reasoning to multi-modal models. Unlike prior efforts, TlDr compresses the length of the reasoning chain without introducing complex data construction and without compromising the original reasoning LLM’s inferential representations.

### v2

# TL;DR: Too Long, Do Re-weighting for Efficient LLM Reasoning Compression

@@ 1 Introduction
Direct approaches aim to addressing the issues of cognitive redundancy and excessive deliberation within reasoning LLMs. Training-free methods CoD ; yao2025activation ; TALE-EP include some that control the internal states of the model during reasoning through prompts or confidence-based techniques to compress the model. Alternatively, the pathway, exemplified by model merging, involves intervening in the parameters of the reasoning LLM to produce relatively concise solutions. Training-based methods primarily focus on sampling and synthesizing relatively concise reasoning paths on specified problem sets through various strategies TokenSkip ; TOPS ; CoT-Valve . These methods involve performing reinforcement learning SimPO ; thinkpruner ; O1-Pruner ; L1 or supervised fine-tuning (SFT) Overthink on reasoning LLMs, enabling the model to learn to generate more concise yet still correct reasoning paths.
@@
We base our approach on an intuitive motivation: when a model is thinking too long, it should re-weight more intuitive reasoning paths to simplify the thinking process. Conversely, when the thinking is too direct, it should incorporate more slow-thinking reasoning chains to encourage deeper contemplation. We propose a dynamic Thinking Length Data Re-Weighting method (TlDr), which dynamically balances the model’s complex reasoning using long CoT and efficient reasoning using short CoT data, enabling the model to eliminate redundant cognitive processes. First, we construct short CoT data for simple problems and long CoT data for complex problems. The model begins with an initial ratio and performs reasoning compression using mixed data. After completing a compression cycle, the model re-evaluates the expected benefits of System-1 CoT data and System-2 CoT data to achieve improved performance. Specifically, and in line with intuition, System-1 CoT data can enhance efficiency, so we use an efficiency metric to measure the expected benefit of System-1 data. System-2 CoT, on the other hand, improves reasoning accuracy, and we use an accuracy metric to measure the benefit of System-2 data in terms of reasoning capability.
Compared to various methods requiring fine-tuning data with different reasoning lengths, our approach enables dynamic ratio learning by utilizing the self-sampled long CoT data and the short CoT data constructed by the original instruct/base model. Through experiments on DeepSeek-Distill-7B/14B, our model achieves excellent compression results on the 7B/14B models, with only a slight decrease in reasoning capability.
@@ 5 Related Work
Despite the strong generalization and reasoning abilities demonstrated by the system 2 reasoning paradigm, the auto-regressive nature of LLMs imposes a significant reasoning burden, limiting their application in agent-based or edge scenarios OS-ATLAS . To address this, various approaches have emerged to improve the reasoning efficiency. These methods can be broadly categorized into two types. One category focuses on building adaptive reasoning-budget. Within this, some training-free methods like CoD CoD and TALE-EP TALE-EP impose budget constraints to control overall reasoning cost. Budget-sensitive models such as L1 L1 , TOPS TOPS , o1 Pruner o1_pruner , and K1.5 team2025kimi add length penalties during the post-training. Some work CoT-Valve ; jiang2025drp ; wang2024explore ; chain-of-reasoning synthesizes diverse-length CoT data, while TOPS TOPS samples budget-sensitive versions using a data model, and C3oT C3oT compresses original LLM output to train jointly on shorter CoTs. Other approaches involve external models or switching mechanisms Reasoning-in-Flux to allocate budgets. Routellm RouteLLM uses multiple routers to find the most suitable reasoning model, while Self-REF Self-REF employs confidence scores to route based on reasoning difficulty. Another category concentrates on building efficient representations. TokenSkip TokenSkip selects data based on token importance for compressed reasoning and more concise thought chains. COCONUT COCONUT explores more efficient reasoning in the latent space. ICoT-KD CCoT and CCoT CCoT attempt to build more efficient reasoning strategies in the hidden space, while Token Assorted combines hidden space and text-based reasoning to balance interpretability and efficiency. Heima Heima extends hidden-space reasoning to multi-modal models. Unlike prior efforts, TlDr compresses the length of the reasoning chain without introducing complex data construction and without compromising the original reasoning LLM’s inferential representations.